OpenStudy (anonymous):

the area of the trapezoid is A=0.5(b1+b2)h find b1 if A=100 h=5 b2=12

OpenStudy (haseeb96):

100=0.5(b1+12)5 100\5=0.5b1+6 20-6 =0.5b1 14\0.5=b1 b1=28 this is the correct answer